3.75 points Required information Sunrise Accounting provides basic tax services and rent-a-controller accounting services.Bluefield has decided to use an ABC costing system and has identified the following detailed information about its cost poolsCost drivers Number of material moves Number of machine hours 500 59,800 3.75 points Required: 1. Calculate Bluefields activ
Answer #1

1) Activity rate

Material handling = 90000/500 = 180 per move

Machine maintenance = 22500/59800 = 0.3763 per MH

Calculate overhead assigned

Product A Product B
Material handling 180*200 = 36000 54000
Machine maintenance 22100*.3763 = 8316 14187
Total Overhead assigned 44316 68187